- 28-06-18Software upgrade core routers
- 27-06-18Onderhoud UPS'en BIT-1
- 26-06-18Periodiek onderhoud UPS'en BIT-2D
- 22-06-18Onderhoud koeling BIT-1
- 22-06-18Periodiek onderhoud UPS'en BIT-2B Feed-B
- 22-06-18Spoedonderhoud UPS-set BIT-2C A-Feed
- 21-06-18Periodiek onderhoud UPS'en BIT-2B Feed-A
- 21-06-18Onderbreking glasvezelverbinding BIT-1 - Equinix AM7
- 21-06-18Spoedonderhoud onderhoud UPS-set BIT-2C A-Feed
- 20-06-18Firewall onderhoud
Maintenance Fiber Ring East BIT-1~BIT-2
On Thursday, September 7th between 00.00h and 07.00h BIT and their fiber installer will perform fiber maintenance on the East single mode fiber ring between BIT-1 and BIT-2.
This maintenance is required to keep in pace with the growing demand for fiber connections. The fiber patch panels on both locations will be replaced, per cable (one per night), by high density patch panels, and the panels will be re-arranged within the cabinet to create more available space for future expansion.
On Wednesday, August 6th between 23:00h and 00:00h preparations will be made for disconnection of the cable, which may result in increased dampening. The actual cut of the fibers will take place after 00:00h.
Because the existing fiber cables need to be cut and re-spliced this maintenance will cause an extensive downtime of several hours for the affected connections.
We kindly ask you to not shutdown the fiber connections, if possible, and only re-route the traffic. This way we can use the light going through the fibers to speed up the patching process after splicing.
If you have one or more redundant fiber connection between BIT-1 and BIT-2 (divided between the East and West ring) these connections will suffer reduced redundancy during this maintenance.
If you have any questions regarding this maintenance, please contact the Customer Care department by phone on +31 (0)318 648 688 or by email on firstname.lastname@example.org.
Update placed at 07/09/2017, 7:00h
All fiberpairs have been reconnected. Unfortunately, we haven't completed testing every pair yet, so it is possible send and receive fibers have been switched. We expect to finish testing and correcting at 7:30h.